A typical light wave has a wavelength of 580 nanometers (580 x 10-9 m).  


 If the speed of light is 3.0 x 108 m/s what is the frequency of the typical light wave?



1
Expert's answer
2022-02-19T04:56:13-0500

The frequency of the light wave

"f=\frac{c}{\lambda}=\frac{3*10^8\:\rm m/s}{580*10^{-9}\rm\: m}=5.17*10^{14}\:\rm Hz"